# Augment then Smooth: Reconciling Differential Privacy with Certified Robustness

### v1

@@ B.1.2 MACER
where $\hat{f}_{\theta}(x)$ denotes the average of softmax probabilties on Gaussian perturbations of $x$, $\hat{f}_{\theta}(x)_{y}$ denotes the softmax probability $\hat{f}_{\theta}(x)$ assigned to the true class $y$, and $\hat{f}_{\theta}(x)_{\hat{y}\neq y}$ denotes the maximum softmax probability for a class that isn’t the true class. This loss term is added to cross entropy of loss from the soft smoothed classifier as a regularization term.
@@ B.2 DPSGD and Robustness
The adversarial robustness of differentially private models has been studied in several prior works. Tursynbek et al. Tursynbek et al. (2021) demonstrated that models trained with DPSGD are sometimes more vulnerable to input perturbations. Boenisch et al. Boenisch et al. (2021) further consolidate this claim with more experiments, and showed that improper choices of hyperparameters can lead to gradient masking. Zhang and Bu Zhang and Bu (2022) find that the success of adversarially training robust models with DPSGD depends greatly on choices of hyperparameters, namely smaller clipping thresholds and learning rates, differing from those that produce the most accurate models. Furthermore, they found that pretraining models helps further mitigate the privacy-robustness-accuracy tradeoff. These works reveal interesting adversarial robustness characteristics of DP models, however, they do not endeavor to improve the robustness of DP models.
DP-ADV Bu et al. (2022a) proposes combining adversarial training with DPSGD. They achieve this by replacing the original example with an adversarially crafted example, obtained with a FGSM or PGD attack. Their work is orthogonal to ours in that they focus on the robustness to adversarial attacks, whereas we focus on certified robustness. Additionally, they do not perform any data augmentation, which has been shown to be effective against adversarial attacks Rebuffi et al. (2021).
There are a few prior works that have studied certified robustness with differential privacy guarantees. Phan et al. Phan et al. (2019) first introduced a framework called Secure-SGD, which aims to achieve both certified robustness and differential privacy simultaneously. They use a PixelDP approach, proposed by Lecuyer et al. Lecuyer et al. (2019), to attain certified robustness and introduced the Heterogeneous Gaussian Mechanism, which involves adding heterogeneous Gaussian noise instead of element-wise Gaussian noise to the gradient. Another work Phan et al. (2020) introduced the StoBatch algorithm to guarantee DP and certified robustness. First, it employs an Autoencoder (AE) Hinton and Zemel (1993) and a functional mechanism (objective perturbation Zhang et al. (2012)) to reconstruct input examples while preserving DP. Subsequently, this reconstructed data is use to train a deep neural network. They implement adversarial training Tramèr et al. (2017) to achieve both robustness and DP for the neural network. Tang et al. Tang et al. (2022) propose transforming input gradients with perturbation during training, and introduced the Multivariate Gaussian Mechanism. This mechanism allows them to achieve the same DP guarantee with less noise added to the gradient. They follow the architecture of denoised smoothing, adding differentially private noise to a pre-trained classifier.
Compared with the existing work, DP-CERT 1) presents a simple and effective augmented training scheme, allowing practitioners to introduce different adversarial training techniques through noising, regularization, and adversarially crafted examples, 2) doesn’t rely on a denoiser at the inference time, reducing inference latency, and 3) can be used for training both randomly initialized or pre-trained networks.
@@ C.1 Dataset Statistics
The CIFAR-10 dataset consists of 60,000 RGB images from 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images of size 32$\times$32$\times$3.
@@ C.2 Code and Implementation
Our code will be released upon publication. We give credit to the original repository for the implementation of SmoothAdv, MACER, stability, and consistency regularization. [^1] https://github.com/jh-jeong/smoothing-consistency For CERTIFY [^2] https://github.com/locuslab/smoothing and local Lipschitz constant [^3] https://github.com/yangarbiter/robust-local-lipschitz evaluations, we use the code provided by the original authors.
@@ D.1 Additional Comparative Study
Figure 3: Approximate certified accuracy comparison on CIFAR10. Baseline results are from Tang et al. (2022).
Figure 2 shows the approximate certified accuracy as the certified radius is increased on MNIST and Fashion-MNIST, corresponding to Table LABEL:tab:certified_accuracy_main in the main text, for different values of the noise $\sigma$ used during training. All variants of DP-CERT greatly outperform the baseline methods which do not specifically encourage robustness. Similar plots for CIFAR10 were shown in Figure LABEL:fig:certified_accuracy_main in the main text. We also compare implementations of DP-CERT to the prior approaches TransDenoiser, SecureSGD, and StoBatch on CIFAR10 in Figure 3. All variants of DP-CERT achieve state-of-the-art certified accuracy on CIFAR10 for differentially private models, with a much smaller pre-trained model compared to Tang et al. (2022).
@@ D.2 Additional Ablation Study
To extend Figure LABEL:fig:ablation from the main text, we show the complete ablation study results in Figures 4 and 5 on MNIST and Fashion-MNIST under $\sigma\in\{0.25,0.5,1.0\}$. Figure 4 shows the effects of adding consistency regularization, or changing the DP clipping method to PSAC. Neither approach makes significant changes to the certified accuracy. Figure 5 shows the effects of changing the multiplicity of augmentations. Again, there is little difference in certified accuracy when using more than two augmentations, so we advocate using the smallest amount which is least expensive computationally. We emphasize that using no augmentations is significantly worse than using two or more – the case without augmentations is simply DPSGD which was compared in Table LABEL:tab:certified_accuracy_main and Figures LABEL:fig:certified_accuracy_main and 2. One of our main conclusions is that adding a small number of Gaussian augmentations to DPSGD is sufficient to greatly improve certified robustness.
Figure 5: Certified accuracy comparison under different numbers of augmentations (2, 4, 6 and 8) on MNIST (top) and Fashion-MNIST (bottom)
@@ D.3 Additional Per-sample Metric Analysis
Figures 6 through 10 show the distributions of our three proposed metrics for interpreting robustness, the input gradient norms, input Hessian spectral norms, and local Lipschitz constants. We use various training methods on MNIST and Fashion-MNIST under $\sigma\in\{0.25,0.5,1.0\}$. Echoing the analysis of RQ1 in Section LABEL:sec:fine_grained_analysis, DPSGD produces a bimodal distribution for the Hessian and gradient norms, while Regular training exhibits a log-normal distribution and smaller tails for large metric values. PSAC shifts the distributions to be closer to those of Regular training by reducing clipping bias. DP-CERT methods, on the other hand, shift the distribution towards smaller metric values, resulting in higher certified accuracy. An exception is DP-Stability, which has significantly higher average gradient and Hessian norms, but without the mode at very high values, and with lower local Lipschitz constants than the other three variants.
(a) Comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 6: Per-sample metric comparisons on MNIST with $\sigma=0.25$
(a) Comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 7: Per-sample metric comparisons on MNIST with $\sigma=1.0$
(a) Comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 8: Per-sample metric comparisons on Fashion-MNIST with $\sigma=0.25$
(a) Comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 9: Per-sample metric comparisons on Fashion-MNIST with $\sigma=0.5$
(a) Comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 10: Per-sample metric comparisons on Fashion-MNIST with $\sigma=1.0$
@@
While our main focus has been on certified robustness, we briefly compare to adversarial robustness against common attacks. Figure 12 shows the adversarial accuracy under a $l_{\infty}$-FGSM attack, with the attack strength in $\{0.0005,0.01,0.1,0.5,1\}$. Consistent with the ranking of the average local Lipschitz constant from Figures 6 through 10, DP-Stability consistently outperforms other approaches, while DP-Gaussian, DP-SmoothAdv, and DP-MACER all achieve similar adversarial accuracy above that of the unprotected baselines.
@@
Finally, we study the metric distributions for individual samples that can or cannot be certified above a threshold $\tau$, similar to Figure LABEL:fig:analysis-mnist-0.5-local-lipschitz in the main text. Figures 13 and 14 respectively show the input gradient norm and input Hessian spectral norm distributions for the baselines and proposed methods, with different threshold values $\tau\in\{0.25,0.5,1.0\}$. In each subfigure the top row shows samples that can be certified at radius larger than $\tau$, and the bottom row shows samples that cannot. We only show the plots for MNIST under $\sigma=0.5$ for brevity. We note that the shapes of the distributions between gradients and Hessians resemble each other closely. The analysis of RQ2 in Section LABEL:sec:fine_grained_analysis also applies here; for example, in Figure 13, the samples with certified radii below the threshold have slightly higher average input gradient norm. As the threshold $\tau$ increases, more examples with higher input gradient norm end up below the certified radius threshold. We find that the samples on which the models are least robust tend to be samples where the gradient and Hessian norms are largest. From this observation we expect that training methods that reduce the prevalence of large gradient and Hessian norms should be the most robust, which is indeed confirmed by DP-Stability which does not have the mode at very large values in Figures 6 through 10, and the best adversarial robustness in Figure 12.

### v3

@@ B.1.3 MACER
where $\hat{f}_{\theta}(x)$ denotes the average of softmax probabilties on Gaussian perturbations of $x$, $\hat{f}_{\theta}(x)_{y}$ denotes the softmax probability $\hat{f}_{\theta}(x)$ assigned to the true class $y$, and $\hat{f}_{\theta}(x)_{\hat{y}\neq y}$ denotes the maximum softmax probability over classes other than the true class. This loss term is added as a regularization term to the cross entropy loss from the soft smoothed classifier.
@@ B.2 DPSGD and Robustness
The adversarial robustness of differentially private models has been studied in several prior works. Tursynbek et al. (2021) demonstrated that models trained with DPSGD are sometimes more vulnerable to input perturbations. Boenisch et al. (2021) further consolidate this claim with more experiments, and showed that improper choices of hyperparameters can lead to gradient masking. Zhang & Bu (2022) found that the success of adversarially training robust models with DPSGD depends greatly on choices of hyperparameters, namely smaller clipping thresholds and learning rates, differing from those that produce the most accurate models. These works reveal interesting adversarial robustness characteristics of DP models, however, they do not endeavor to improve the robustness of DP models.
Bu et al. (2022a) propose combining adversarial training with DPSGD. They achieve this by replacing the original example with an adversarially crafted example, obtained with a FGSM or PGD attack. Their work is orthogonal to ours in that they focus on the robustness to adversarial attacks, whereas we focus on certified robustness. Additionally, they do not perform any data augmentation, which has been shown to be effective against adversarial attacks (Rebuffi et al., 2021).
Compared with existing works in DP and CR (Phan et al., 2019; 2020; Tang et al., 2022) reviewed in Section 2, DP-CERT 1) presents a simple and effective training scheme with augmentations, allowing practitioners to introduce different adversarial training techniques through noising, regularization, and adversarially crafted examples, 2) does not rely on a denoiser at inference time, reducing inference latency, and 3) can be used for training both randomly initialized or pre-trained networks.
@@ C.1 Datasets
The CIFAR10 dataset consists of 60,000 RGB images from 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images of size 32$\times$32$\times$3.
@@ C.2 Code and Implementation
Our implementation of DP-CERT in code is provided as supplementary material. We give credit to the original repository for the implementation of SmoothAdv, MACER, Stability, and Consistency regularization. [^1] https://github.com/jh-jeong/smoothing-consistency For CERTIFY [^2] https://github.com/locuslab/smoothing and computing local Lipschitz constants [^3] https://github.com/yangarbiter/robust-local-lipschitz, we use the code provided by the original authors. All experiments were conducted on a cluster of 8 Nvidia V100 GPUs. All the training and inference procedures are implemented based on Pytorch v1.13.0 (Paszke et al., 2019) and Opacus v1.3.0 (Yousefpour et al., 2021), and our code is provided at github.com/layer6ai-labs/dp-cert.
C.3 Additional Implementation and Hyperparameters
We set the learning rate as 0.001 and train the models for 10 epochs. The rest of the hyperparameters are the same as used by Bu et al. (2022b). For evaluation, we use CERTIFY with parameters $n=10,000$, $n_{0}=100$, and $\alpha=0.001$, following previous work (Cohen et al., 2019; Salman et al., 2019). Our implementations of DP-SmoothAdv, DP-Stability and DP-MACER are adapted from the original codebases for those regularization approaches (Salman et al., 2020; Li et al., 2019; Zhai et al., 2020), and we use the default hyperparameters reported in the original papers. We set the number of augmentations $K$ to 2 for MNIST and Fashion-MNIST, and 1 for CIFAR10, as they bring a better trade-off between certified accuracy and efficiency (see Figure LABEL:fig:ablation).
@@ Appendix D Additional Results
In this section, we show the complete results and figures for our comparative study, ablations, and analysis of fine-grained metrics.
@@ D.1 Additional Comparative Study
Figure 2 complements Figure LABEL:fig:certified_accuracy_main from the main text by comparing the approximate certified accuracy between DP-CERT and baselines on MNIST, and Fashion-MNIST. The results confirm that Regular and DP training methods that do not specifically consider robustness as a criteria give poor CR results, but all instantiations of the DP-CERT framework do much better. We are unable to show further results similar to Figure LABEL:fig:certified_accuracy_transdenoiser using the baselines SecureSGD, StoBatch, and TransDenoiser because their implementations are highly customized and not compatible with our codebase based on Pytorch and Opacus. The baseline results in Figure LABEL:fig:certified_accuracy_transdenoiser are directly copied from (Tang et al., 2022) which provided limited experimental settings for us to match.
Figure 4: Certified accuracy comparison under different numbers of augmentations (0, 2, 4, 6 and 8) on MNIST (top) and Fashion-MNIST (bottom). The line marked 0 indicates replacing the original datapoint with an augmentation.
@@ D.2 Additional Ablation Study
To extend Figure LABEL:fig:ablation from the main text, we complete the ablation study in Figures 3 and 4 on MNIST and FashionMNIST under $\sigma\in\{0.25,0.5,1.0\}$. Figure 3 shows the effects of adding Consistency regularization, or changing the DP clipping method to PSAC. Neither modification leads to significant changes to the certified accuracy, so we recommend using the simplest combination of standard DPSGD clipping without additional regularization when training models from scratch. Figure 4 shows the effects of changing the multiplicity of augmentations $K$. Since additional augmentations greatly increase computation time, we must navigate the tradeoff between efficiency and performance. We see that it is beneficial to use some number of augmentations or replace the original datapoint with a single augmentation (marked line 0), but beyond $K=2$ there is little further gain in certified accuracy. We note that the case without augmentations at all, namely DPSGD, achieves much worse CR as presented in Table LABEL:tab:certified_accuracy_main and Figures LABEL:fig:certified_accuracy_main and 2. One of our main conclusions is that adding a small number of Gaussian augmentations to DPSGD is sufficient to improve certified robustness at a low cost to efficiency.
Table 2 and 3 show additional hyperparameter studies on learning rate $\lambda$ and clipping bound $C$ for DP-Gaussian under $\sigma=0.25$. Not only does this show that our models were well-tuned for clean accuracy, but that there is no dramatic drop in performance for nearby parameter values.
Table 2: Effect of Learning Rate $\lambda$ on Clean Accuracy (%)
Dataset Learning Rate $\lambda$ Clean Accuracy (%)
MNIST 0.25 97.71
0.5 98.13
1 97.78
Fashion-MNIST 2 83.47
4 84.76
8 83.96
CIFAR10 0.0005 84.11
0.001 87.61
0.002 85.53
Table 3: Effect of Clipping Norm $C$ on Clean Accuracy (%)
Clipping Norm $C$ Clean Accuracy (%)
MNIST Fashion-MNIST CIFAR10
0.01 94.89 75.60 83.79
0.05 97.75 83.42 83.99
0.1 98.13 84.76 87.61
0.2 97.72 84.00 84.19
0.5 96.47 81.46 84.24
@@ D.3 Additional Per-sample Metric Analysis
Figures 5 through 9 complete our presentation of the three proposed metrics for interpreting robustness, the input gradient norms, input Hessian spectral norms, and local Lipschitz constants, over three datasets and values of $\sigma$ (see also Figure LABEL:fig:analysis-mnist-0.5 in the main text). We use various training methods on MNIST and Fashion-MNIST under $\sigma\in\{0.25,0.5,1.0\}$. Echoing the analysis of RQ1 in Section LABEL:sec:fine_grained_analysis, DPSGD produces a bimodal distribution for the Hessian and gradient norms, while Regular training exhibits a log-normal distribution and smaller tails for large metric values. PSAC shifts the distributions to be closer to those of Regular training by reducing clipping bias. DP-CERT methods, on the other hand, shift the distribution towards smaller metric values, resulting in higher certified accuracy. An exception is DP-Stability, which has significantly higher average gradient and Hessian norms, but without the mode at very high values, and with lower local Lipschitz constants than the other three variants.
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 5: Per-sample metric comparison, MNIST, $\sigma=0.25$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 6: Per-sample metric comparison, MNIST, $\sigma=1.0$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 7: Per-sample metric comparison, Fashion-MNIST, $\sigma=0.25$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 8: Per-sample metric comparison, Fashion-MNIST, $\sigma=0.5$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 9: Per-sample metric comparison, Fashion-MNIST, $\sigma=1.0$
@@
While our main focus has been on certified robustness, we briefly examine adversarial robustness. Figure 11 shows the adversarial accuracy under a $l_{\infty}$-FGSM attack, with the attack strength in $\{0.0005,0.01,0.1,0.5,1\}$. Consistent with the ranking of the average local Lipschitz constant from Figures 5 through 9, DP-Stability consistently outperforms other approaches, while DP-Gaussian, DP-SmoothAdv, and DP-MACER all achieve similar adversarial accuracy above that of the unprotected baselines.
Figure 12: Distributions of log local Lipschitz constants among baselines and proposed methods, MNIST, $\sigma=0.5$. In each subfigure, data points are classified into ones with certified radius above the threshold $\tau$ (top row), and below the threshold (bottom row). We display $\tau\in\{0.5,1.0,1.5\}$.
@@
Finally, as shown in Figure 12, we select three certified radius thresholds $\tau\in\{0.5,1.0,1.5\}$, and separately plot the log of local Lipschitz constants of examples above and below $\tau$ on the top and bottom rows of the subfigures respectively. First, the examples with certified radii below the threshold have higher average local Lipschitz constant. Second, as we increase the threshold $\tau$, more examples with higher local Lipschitz constant end up below the certified radius threshold. We do a similar analysis for the other two metrics in Figures 13 and 14, but since the local Lipschitz constant is derived using a PGD attack, it naturally correlates with the robustness to adversarial examples better. For further comparison, we present the FGSM accuracy on MNIST and Fashion-MNIST under attack strengths in $\{0.0005,0.01,0.1,0.5,1\}$ in Figure 11. Consistent with the ranking of the average local Lipschitz constant, DP-Stability consistently outperforms other approaches, while DP-Gaussian, DP-SmoothAdv and DP-MACER achieve similar adversarial accuracy over different attack margins.